応用解析学特論 2 単位 (選択)

Advanced applied analysis  岡本邦也 · 講師/ 知的力学システム工学専攻建設創造システム工学コース社会システム工学講座

【授業目的】数理科学の基礎的道具である関数解析的手法について学ぶ．【授業概要】無限次元空間における線形代数にあたる関数解析学の基礎理論を講

義し，微分方程式で記述される諸現象に対する関数解析的アプローチを紹介する．本科目は，工業に関する科目である．

【授業形式】講義【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】『数理解析特論』 (0.5) 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】現代解析学の理論が適用でき，且つその有用性を認識できる．【授業計画】

1. 微分方程式とその解

2. 行列の指数関数 1

3. 解の一意性

4. 解の存在

5. 固有値と固有空間

6. 射影分解

7. 行列の指数関数 2

8. 一般固有値問題

9. Dunford 積分

10. 行列の正則関数

11. 解曲線と安定性

12. 解の安定性

13. リヤプノフの方法

14. 非線型の場合

15. 線型近似

16. 総括【成績評価基準】授業中に課すレポートで評価する．【教科書】 [教科書] 【参考書】『新微分方程式対話』笠原晧司著，日本評論社【WEB 頁】 http://math9.pm.tokushima-u.ac.jp/lecture/ 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216571 【対象学生】 [対象学生] 【連絡先】

⇒ 岡本(A棟212室, TEL/FAX: 088-656-9441, E-mail: okamoto@pm.tokushim a-u.ac.jp) (オフィスアワー: 【WEB頁】のHPを参照のこと)

Advanced applied analysis 2 units (selection)

Kuniya Okamoto · Associate Professor / Planning and Design Systems Engineering for Infrastructures, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

Target⟩ To learn the functional analytic methods which are basic tools for

mathematical sciences. Outline⟩ This subject provides basic theory of functional analysis which is

considered as linear algebra in infinite-dimensional spaces. Functional analytic approaches to phenomena described by differential equations are introduced.

Style⟩ Lecture �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Topics of Analysis for Mathematical Science”(0.5) Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ To apply the theory of modern analysis and recognize its significant role.

Schedule⟩

1. Differential equations and their solutions

2. Exponential of matrices 1

3. Uniqueness of solutions

4. Existence of solutions

5. Eigenvalues and eigenspaces

6. Projective representation

7. Exponential of matrices 2

8. Generalized eigenvalue problems

9. Dunford integrals

10. Holomorphic functions of matrices

11. Solution curve and stability

12. Stability of solutions

13. Ljapunov’s method

14. Nonlinear case

15. Linear approximations

16. Summary Evaluation Criteria⟩ Evaluation by the report. Textbook⟩ [教科書] Reference⟩ 『新微分方程式対話』笠原晧司著，日本評論社 Webpage⟩ http://math9.pm.tokushima-u.ac.jp/lecture/ Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216571 Student⟩ [対象学生] Contact⟩

⇒ 岡本(A棟212室, TEL/FAX: 088-656-9441, E-mail: okamoto@pm.tokushim a-u.ac.jp) (Office Hour: 【WEB頁】のHPを参照のこと)